OpenStudy (anonymous):

Simplify (7g2 – 4g + 9) + (6g2 + 8g – 1)

OpenStudy (anonymous):

i.e. add?

OpenStudy (anonymous):

\(7g^2+6g^2=13g^2\) right?

OpenStudy (anonymous):

\(8g-4g=4g\) and \(9-1=8\) giving you a final answer of \[13g^2+4g+8\]

OpenStudy (anonymous):

I just got that.. Wow. It seems common sense doesn't come naturally to me. Thank you :P

OpenStudy (anonymous):

yw